How many calories are coming from protein only in the Quarter pounder?

Calorie Content of Macronutrients 

  • Fat: 9 calories per gram 

  • Protein: 4 calories per gram 

  • Carbohydrates: 4 calories per gram 

  • Alcohol: 7 calories per gram 

Please answer the following questions. You may refer to the internet and your textbook for assistance. Show all work. 

1) How many calories are coming from protein only in the Quarter pounder? What about carbs and fat separately? 
Answer #1

A) Total calories from protein in Quarter Pounder (with cheese) = 264 calories

Calculation based on the nutritional values from mcdonalds.com

Weight of one Quarter Pounder = 220 grams

Protein content in 100 grams of Quarter Pounder = 30 grams

Protein content in 220 grams or one Quarter Pounder = 220 * 30/100 = 220 * 30/100 = 66 grams

As per the calorie content of macronutrients given in the question, one gram protein has 4 calories

Therefore, total calories from protein in a Quarter Pounder = 66 * 4 = calories = 264 calories

B) Total calories from carbs and fat separately in a Quarter Pounder

Total calories from fat in a Quarter Pounder = 495 calories

Total calories from carbs in a Quarter Pounder = 369.6 calories

Calculation based on the nutritional values from mcdonalds.com

Weight of one Quarter Pounder = 220 grams

Total fat content in 100 grams of Quarter Pounder = 25  grams

Total fat content in 220  grams or one Quarter Pounder = 220 * 25/100 = 220 * 0.25 = 55 grams

As per the calorie content of macronutrients given in the question, one gram fat has 9 calories

Therefore, total calories from fat in a Quarter Pounder = 55 * 9 = 495 calories

Carb content in 100 grams of Quarter Pounder = 42 grams

Carb content in 220  grams or one Quarter Pounder = 220 * 42/100 = 220 * 0.42 = 92.4 grams

As per the calorie content of macronutrients given in the question, one gram of carbohydrates has 4 calories

Therefore, total calories from carbs in a Quarter Pounder = 92.4 * 4 = 369.6 calories